# Audio-Visual Person Verification based on Recursive Fusion of Joint Cross-Attention

### v2

@@ IV Results and Discussion
A) Dataset and Evaluation Metrics: We have evaluated the proposed model on the Voxceleb1 dataset [25], obtained from YouTube videos under challenging environments. The dataset consists of 1,48,642 video clips, captured from 1,251 speakers, out of which 55% of the speakers are male. Each video clip has a duration of 4 to 145 seconds, and the speakers are chosen to cover a diverse range of ethnicities, accents, professions, and ages. For our experiments, we have divided the voxceleb1 development set, which has 1211 speakers into training and validation sets. The training and validation splits were randomly selected as 1150 and 61 speakers respectively and reported our results on both the validation split and Vox1-O (Voxceleb1 original) test set for performance evaluation. It is worth mentioning that the models are trained only on the Voxceleb1 dataset.
@@
B) Ablation Study: The results are reported based on the average of three runs for statistical stability. To obtain the audio and visual feature vectors, we have used Resnet-18 [28] for visual modality and ECAPA-TDNN [29] for audio modality similar to that of [4] in order to have a fair comparison. We have conducted a series of experiments to analyze the performance of the proposed approach by comparing it with some of the widely used fusion strategies as shown in Table 1. First, we have implemented a simple score-level fusion, where scores are obtained from the individual modalities and then fused to verify the identity of a person. We also explored another widely used fusion strategy of feature concatenation, where the features of audio and visual modalities are concatenated and used to obtain the final score. We can observe that the fusion performance has been improved over simple score-level fusion. To analyze the impact of the proposed approach, we further compared it with some of the widely explored attention mechanisms. By employing a self-attention mechanism on the concatenated features of individual modalities, the fusion performance has been further improved by leveraging the intra-modal relationships. We also explored inter-modal relationships across the modalities using cross-attention and found further improvement in fusion performance over prior fusion strategies. Now, we explored joint cross-attention, where joint feature representation is deployed in the cross-attention framework to simultaneously capture both intra- and inter-modal relationships. Finally, we have introduced the proposed recursive fusion of the attended features of individual modalities, and we can observe that recursive fusion helps in obtaining more refined feature representations and achieves the best performance among all the fusion strategies.
To further analyze the contribution of the individual components of the proposed approach, we conducted another series of experiments with BLSTMs and varying the number of iterations of the recursive mechanism. First, we performed two experiments without recursion, one with BLSTM and another without BLSTM to understand the contribution of BLSTM to the proposed approach. Table II shows the results of the experiments on the validation set of the Voxceleb1 dataset. We can observe that BLSTMs improve the fusion performance by taking advantage of the temporal modeling of the A-V feature representations. To understand the impact of the recursive fusion, we have performed experiments by varying the number of iterations and obtained the best performance at 3 iterations. Beyond that, we observe a decline in the fusion performance, which can be attributed to the fact that though recursion helps in improving fusion performance initially, it results in overfitting with more iterations, resulting in a performance decline on the validation set. We also performed experiments with multiple iterations but without BLSTMs and found a similar trend of improvement in fusion performance.
@@
Method EER normal-$\downarrow$ minDCF normal-$\downarrow$
@@
Method EER normal-$\downarrow$ minDCF normal-$\downarrow$
@@
Method EER normal-$\downarrow$ minDCF normal-$\downarrow$ EER normal-$\downarrow$ minDCF normal-$\downarrow$
@@
C) Comparison to state-of-the-art: We have used the Voxceleb1 dataset to validate the proposed approach and compared it with the relevant state-of-the-art methods using A-V fusion for person verification. To have a fair comparison, we have re-implemented the state-of-the-art methods using our experimental setup due to the differences in their experimental protocols. Table III shows the comparison of the proposed approach to state-of-the-art methods as well as individual modalities on both validation split of Voxceleb1 and Vox1-O datasets. First, we have conducted experiments with the individual modalities and found that audio performs relatively better than visual modality.
Tao et al [4] explored the complementary relationships across the audio and visual modalities as supervisory information to mitigate the impact of noisy labels. We have used the clean samples obtained using [4] to train our models as well as other methods [17, 10, 21]. Most of the existing approaches [17, 10, 4] rely on a single image (randomly chosen), thereby focusing on the inter-modal relationships, failing to retain the intra-modal relationships. Both of these approaches [17] and [10] are able to improve the fusion performance better than that of [4], where [17] employs shared representation space using a shared classifier and [10] explored gating mechanism. Mocanu et al [21] also focused on capturing the inter-modal relationships based on cross-correlation across the modalities. Unlike these approaches, we have focused on capturing both intra- and inter-modal relationships simultaneously by introducing the joint feature representation in the cross-attention framework in a recursive fashion. Since the proposed RJCA model leverages both intra- and inter-modal relationships effectively, we can observe that the fusion performance has been improved over other state-of-the-art methods.
@@ V CONCLUSION
In this paper, we have presented a novel approach of recursive joint cross-attentional fusion for person verification by effectively exploiting both inter- and intra-modal relationships across audio and visual modalities. Specifically, we have explored the joint feature representation in the cross-attentional framework recursively to obtain more refined A-V feature representations. To further enhance the intra-modal relationships of A-V feature representations, we have used BLSTMs on the refined A-V feature representations obtained from the proposed RJCA model. We showed that the fusion performance can be improved by effectively capturing the intra- and inter-modal relationships. The performance of the proposed approach can be further enhanced by training with the large-scale Voxceleb2 dataset as it can improve the generalization ability of the proposed approach.

### v3

@@ IV Results and Discussion
A) Dataset and Evaluation Metrics: We have evaluated the proposed model on the Voxceleb1 dataset [27], obtained from YouTube videos under challenging environments. The dataset consists of 1,48,642 video clips, captured from 1251 speakers, of which 55% of the speakers are male. Each video clip has a duration of 4 to 145 seconds, and the speakers are chosen to cover a diverse range of ethnicities, accents, professions, and ages. For our experiments, we have divided the voxceleb1 development set, which has 1211 speakers into training and validation sets. The training and validation splits were randomly selected as 1150 and 61 speakers, respectively, and our results are reported on both the validation split and Vox1-O (Voxceleb1 original) test set for performance evaluation. It is worth mentioning that the models are trained only on the Voxceleb1 dataset.
@@
B) Ablation Study: The results are reported based on the average of three runs for statistical stability. To obtain the audio and visual feature vectors, we have used Resnet-18 [30] for visual modality and ECAPA-TDNN [31] for audio modality similar to that of [4] in order to have a fair comparison. We have conducted a series of experiments to analyze the performance of the proposed approach by comparing it with some of the widely used fusion strategies shown in Table 1. First, we have implemented a simple score-level fusion, where scores are obtained from individual modalities and then fused to verify the identity of a person. We also explored another widely used fusion strategy of feature concatenation, where the features of the audio and visual modalities are concatenated and used to obtain the final score. We can observe that the fusion performance has been improved over simple score-level fusion. To analyze the impact of the proposed approach, we further compared it with some of the widely explored attention mechanisms. By employing a self-attention mechanism on the concatenated features of individual modalities, the fusion performance has been further improved by leveraging the intra-modal relationships. We also explored inter-modal relationships across the modalities using CA and found further improvement in fusion performance over prior fusion strategies. Now, we employed joint cross-attention, where joint feature representation is integrated into the CA framework to simultaneously capture both intra- and inter-modal relationships. Finally, we have introduced the proposed recursive fusion with joint CA to obtain the attended features of individual modalities, and we can observe that recursive fusion helps in obtaining more refined feature representations and achieves the best performance among all the fusion strategies.
To further analyze the contribution of the individual components of the proposed approach, we conducted another series of experiments with BLSTMs and varying the number of iterations of the recursive mechanism. First, we performed two experiments without recursion, one with BLSTM and another without BLSTM to understand the contribution of BLSTM to the proposed approach. Table II shows the results of the experiments on the validation set of the Voxceleb1 dataset. We can observe that BLSTMs improve the fusion performance by taking advantage of the temporal modeling of the A-V feature representations. To understand the impact of the recursive fusion, we performed experiments by varying the number of iterations and obtained the best performance at 3 iterations. Beyond that, we observe a decline in the fusion performance, which can be attributed to the fact that though recursion helps in improving fusion performance initially, it results in overfitting with more iterations, resulting in a performance decline on the validation set.
@@
Method EER $\downarrow$ minDCF $\downarrow$
@@
Method EER $\downarrow$ minDCF $\downarrow$
@@
Method EER $\downarrow$ minDCF $\downarrow$ EER $\downarrow$ minDCF $\downarrow$
@@
C) Comparison to state-of-the-art: We have used the Voxceleb1 dataset to validate the proposed approach and compared it with the relevant state-of-the-art methods using A-V fusion for person verification. To have a fair comparison, we have reimplemented the state-of-the-art methods using our experimental setup due to the differences in their experimental protocols. Table III shows the comparison of the proposed approach with state-of-the-art methods and individual modalities on both the validation split of the Voxceleb1 and Vox1-O partitions. First, we have conducted experiments with the individual modalities and found that audio performs relatively better than visual modality.
Tao et al. [4] explored complementary relationships between the audio and visual modalities as supervisory information to mitigate the impact of noisy labels. We have used the clean samples obtained using [4] to train our models as well as other methods [17, 10, 22, 23]. Most of the existing approaches [17, 10, 4] rely on a single image (randomly chosen), thereby focusing on the inter-modal relationships, failing to retain the intra-modal relationships. Both of these approaches [17] and [10] are able to improve the fusion performance better than that of [4], where [17] uses a shared representation space using a shared classifier and [10] explored the gating mechanism. Mocanu et al. [22] also focused on capturing the inter-modal relationships based on cross-correlation across the modalities. Unlike these approaches, Praveen et al. [23] showed improvement by capturing both intra- and inter-modal relationships simultaneously using joint feature representation in the CA framework. We have further improved the fusion performance by introducing recursive fusion to obtain more refined feature representations. Since the proposed RJCA model leverages both intra- and inter-modal relationships effectively, we can observe that the fusion performance has been improved over other state-of-the-art methods.
@@ V CONCLUSION
In this paper, we have presented a novel approach of recursive joint CA fusion for person verification by effectively exploiting both inter- and intra-modal relationships across audio and visual modalities. Specifically, we have explored the joint feature representation in the CA framework recursively to obtain more refined A-V feature representations. To further enhance the intra-modal relationships of A-V feature representations, we have used BLSTMs on the refined A-V feature representations obtained from the proposed RJCA model. We showed that the fusion performance can be improved by effectively capturing the intra- and inter-modal relationships. The performance of the proposed approach can be further enhanced by training with the large-scale Voxceleb2 dataset as it can improve the generalization ability of the proposed approach.